OpenStudy (anonymous):

How to solve x^2 – 10x = –13?

OpenStudy (anonymous):

The choices are 5 ± 2√3 5 ± √38 5 ± 2√2 5 ± 3√2 but i dont just want an answer i need help on how to do the problem.

OpenStudy (anonymous):

complete the square by writing \[(x-5)^2=-13+25=12\] then take the square root

OpenStudy (jiteshmeghwal9):

\(\LARGE{x^2-10x+13=0}\) Here, a=1, b=-10, c=13. use the quadratic formula\[\LARGE{x={-b \pm \sqrt{b^2-4ac} \over 2a}}\] \[\LARGE{x={-(-10) \pm \sqrt{(-10)^2-4(1)(13)}\over 2(1)}}\]

OpenStudy (abb0t):

Add 13 to both sides and set it equal to zero. use quadratic formula to solve: \[\frac{ - b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (abb0t):

Plug it in where \[ax^2 + bx + c = 0\]

OpenStudy (jiteshmeghwal9):

Solve this equation for 'x' :)

OpenStudy (anonymous):

Ok i think i have a better understanding.

OpenStudy (anonymous):

When I work out whats in the radical i get an irational number what do i do?

OpenStudy (anonymous):

Factor the number in the radical. If you have any factor repeated twice, pull it out.

OpenStudy (anonymous):

What i got was 4*1=4 4*13=52 -10^2=100 100-52=48 then i forget how to factor the number from there